OpenStudy (anonymous):

A particle moves in the xy-plane along the path given by y=5x^2+3x+10. Thex-component of the velocity, V(x), is a constant 4 m/s and at x=x(0)=6m. What is x and y as a function of t? Find d^2y/dt and d^2x/dt.

OpenStudy (anonymous):

dx/dt = 4 so dx = 4dt integrate x = 4t + C since we know x(0) = 4*0 + C =6 we get C =6 and we know : x = 4t + 6 now you can find y(t) and hence the derivatives..

OpenStudy (anonymous):

Ugh... Thank you. I'm a math major, and this is basic Calc. Studying so long that I'm running on E. Lol.

OpenStudy (anonymous):

you know how to do the rest ?

OpenStudy (anonymous):

Yeah, yeah... basic integrals and derivatives. Just haven't seen a dy/dx problem in a while. Took my entire math rotation before taking the physics. Now i'm rusty on my basic calc.